Circular No. 7824

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2002 BM_{26}

M. C. Nolan and E. S. Howell, National Astronomy and Ionosphere Center; C. Magri and B. Beeney, University of Maine; D. B. Campbell, Cornell University; L. A. M. Benner, S. J. Ostro, and J. D. Giorgini, Jet Propulsion Laboratory; and J.-L. Margot, California Institute of Technology, write: “Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on Feb. 9 and 10 show that 2002 BM_{26} (cf. MPEC 2002-C14) is a binary system. Preliminary estimates of average diameters, based on range extents at 15-m and 30-m resolution, are 600 and 100 m. On Feb. 10.05 UT, the two objects were measured to be at the same distance, to within 100 m. The rotation period of the primary is ∼ 2.7 hr. The orbital period, not yet well constrained, is not more than three days.”

1997 CQ_{29}

K. Noll and D. Stephens, Space Telescope Science Institute; report for their team (W. Grundy, J. Spencer, R. Millis, M. Buie, Lowell Observatory; D. Cruikshank, Ames Research Center, NASA; S. Tegler, Northern Arizona University; W. Romanishin, University of Oklahoma) the identification of a possible binary transneptunian object (TNO): “During 2001 Nov. 17.281– 17.307 UT, 1997 CQ_{29} (MPEC 1997-J02; MPO 10838) was observed with the WFPC2 camera on the Hubble Space Telescope in three broadband filters approximating V, R, and I, with two exposures in each filter. All six images of 1997 CQ_{29} are extended by 0′′.17 ± 0′′.03 along an axis in p.a. 17^{o}. The orientation was approximately diagonal on the WF3 chip. The telescope was tracking the apparent motion of the TNO during the observations in a direction nearly perpendicular to the elongation. The centroid of the object did not move measurably between the first and last observation, so tracking errors are not the cause of the elongation. Only one other TNO, out of some 60 TNOs observed by us with HST, has a similar appearance (the binary 1998 WW_{31}; cf. IAUC 7610). With 1997 CQ_{29} at a geocentric distance of 41.85 AU, the long axis of its image corresponds to a span of 5200 km. With a Johnson V magnitude of 23.42 ± 0.08, we conclude that 1997 CQ_{29} must be a binary, rather than a single elongated object. We also measure Cousins magnitudes R = 22.64 ± 0.06 and I = 22.14 ± 0.06, giving this object typical TNO colors.”

SUPERNOVA 2002au IN UGC 5100 SN 2002au was found on the KAIT images (IAUC 7823) by P. Martin.

2002 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green